# GroverGPT: A Large Language Model with 8 Billion Parameters for Quantum Searching

### v1

@@ I Introduction
As a first step toward understanding these simulation capabilities, we analyze our approach using quantum-specific metrics – search accuracy ($\alpha$), fidelity ($\epsilon$), and marked fidelity ($\epsilon^{k}$). Our preliminary results show promising performance in simulating quantum search on moderate-sized systems, with encouraging signs of generalization from training on small systems ($3\sim 6$ qubits) to somewhat larger quantum registers (e.g., $20$ qubits). While these initial findings suggest interesting possibilities for classical simulation of quantum algorithms, they also reveal important limitations that appear to be fundamental rather than technical. This exploration provides early insights into the challenges and opportunities at the boundary between classical and quantum computation. This initial investigation contributes to quantum computing research through four key aspects:
@@
  • Initial empirical observations about how classical systems might learn and generalize quantum principles without explicit quantum state representation. We explore a potential approach to quantum algorithm simulation that aims to balance fidelity with computational efficiency.
@@ II GroverGPT
Figure 3: Performance evaluation and generalization capability of GroverGPT. (a) Distribution of the pre-training dataset comprising $97$K quantum search examples across different qubit sizes from $3$ to $20$. (b) Training loss curves for different GroverGPT variants, showing convergence behavior during pre-training on 3-6 qubit datasets. (c) Comparative accuracy analysis of GPT-4o, GroverGPT, and its variants with QASM and conversation components on $6$-qubit (left) and $10$-qubit (right) test sets across different training qubit ranges. (d) Fidelity $\epsilon$ comparison between models on $6$-qubit (left) and $10$-qubit (right) test sets, demonstrating the error reduction as training qubit count increases. (e) Scalability assessment of GroverGPT trained on $3\sim 6$ qubits, showing accuracy (red line) and Fidelity (blue bars) when tested on larger systems ranging from $6$ to $20$ qubits, highlighting the LLM’s generalization capabilities beyond its training domain.
@@ II.3 Scalability and Generalization Capability of GroverGPT
In this section, we examine GroverGPT’s remarkable ability to generalize quantum search capabilities from training on small-scale systems to significantly larger quantum systems. GroverGPT exhibits remarkable generalization capabilities across different qubit scales. Figure 3e comprehensively analyzes the model’s performance when trained on $3\sim 6$ qubits and tested on increasingly larger systems, ranging from $6$ to $20$ qubits. The results show that GroverGPT maintains accuracy above $95\%$ for systems up to $8$ qubits, with only a gradual decline to approximately $95\%$ for systems between $9$ and $20$ qubits. This robust generalization ability is particularly noteworthy given that the model was trained only on smaller systems ($3\sim 6$ qubits), suggesting its capacity to extrapolate quantum search principles to substantially larger quantum systems. The parallel trends between accuracy and Fidelity in Figure 3e suggest that the model’s performance degradation at larger qubit counts is gradual and predictable, indicating a systematic rather than catastrophic breakdown of simulation capability. This behavior aligns with theoretical expectations about the scalability challenges in quantum simulation and provides valuable insights into the practical limits of classical models in capturing quantum phenomena.
@@
Through detailed fidelity analysis in this section, we investigate whether GroverGPT truly learns quantum search rather than simply approximating classical search strategies. The Fidelity analysis in Figure 3d provides strong evidence that GroverGPT learns genuine quantum search rather than classical approximations. The consistently low Fidelity values (below $0.005$ for $6$ qubits and approaching zero for systems trained on $5$ or more qubits) indicate that GroverGPT accurately captures the quantum state amplitudes characteristic of Grover’s algorithm. This is in stark contrast to GPT-4o openai2024gpt4ocard ’s performance, which shows a consistently high Fidelity (approximately $0.011$), suggesting it fails to capture the subtle quantum features of the search process. The convergence of Fidelity to near-zero values with increased training qubit count is particularly significant, as it indicates that GroverGPT successfully learns the interference patterns and amplitude amplification mechanisms that are quintessential to quantum search. This distinction is further reinforced by the symmetric behavior observed in both $6$-qubit and $10$-qubit test cases, suggesting that the model has internalized fundamental quantum principles rather than memorizing specific problem instances.
@@ III.3 Evaluation Metrics
  • Fidelity: Quantifies the overall quantum state reproduction accuracy through:
@@
  • Marked Fidelity ($\epsilon^{k}$): Specifically evaluates the accuracy of marked state probability predictions:
@@
These metrics provide complementary perspectives on GroverGPT’s performance: $\alpha$ assesses the model’s ability to identify correct search solutions, $\epsilon$ evaluates the overall quantum state reproduction fidelity, and $\epsilon^{k}$ focuses specifically on the accuracy of marked state predictions. Together, they comprehensively evaluate the model’s practical search capabilities and theoretical quantum properties.
@@ Acknowledgment
We thank Yangrui Hu, Zhiding Liang and Suqiong Zeng for discussion and help. MC, JC, and JL are supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, and the PQI Community Collaboration Awards. PL and TC are supported in part by Cisco Faculty Award and UNC SDSS Seed Grant.
@@ IV.5 GroverGPT Fine-Tuning Details
where $N$ is the number of training samples, $T_{i}$ is the length of the target sequence for the $i$-th sample, $y_{i,t}$ is the ground truth token, $y_{i,<t}$ represents the preceding tokens, $x_{i}$ is the input prompt, and $P_{\theta}(y_{i,t}|y_{i,<t},x_{i})$ is the model’s predicted probability.
This formulation ensures that the model learns to predict target tokens accurately based on the provided context and previously predicted tokens.

### v3

@@ I Introduction
As a first step toward understanding these simulation capabilities, we analyze our approach using quantum-specific metrics – search accuracy ($\alpha$), infidelity ($\epsilon$), and marked infidelity ($\epsilon^{k}$). Our preliminary results show promising performance in simulating quantum search on moderate-sized systems, with encouraging signs of generalization from training on small systems ($3\sim 6$ qubits) to somewhat larger quantum registers (e.g., $20$ qubits). While these initial findings suggest interesting possibilities for classical simulation of quantum algorithms, they also reveal important limitations that appear to be fundamental rather than technical. This exploration provides early insights into the challenges and opportunities at the boundary between classical and quantum computation. This initial investigation contributes to quantum computing research through four key aspects:
@@
  • Initial empirical observations about how classical systems might learn and generalize quantum principles without explicit quantum state representation. We explore a potential approach to quantum algorithm simulation that aims to balance infidelity with computational efficiency.
@@
  • The model’s robustness is comprehensively evaluated (Section II.6), highlighting the impact of training strategies and dataset diversity on performance consistency.
@@ II GroverGPT
Figure 3: Performance evaluation and generalization capability of GroverGPT. (a) Distribution of the pre-training dataset comprising $97$K quantum search examples across different qubit sizes from $3$ to $20$. (b) Training loss curves for different GroverGPT variants, showing convergence behavior during pre-training on 3-6 qubit datasets. (c) Comparative accuracy analysis of GPT-4o, various open-source large models, GroverGPT, and its variants with QASM and conversation components on $6$-qubit (left) and $10$-qubit (right) test sets across different training qubit ranges. (d) Infidelity $\epsilon$ comparison between models on $6$-qubit (left) and $10$-qubit (right) test sets, demonstrating the error reduction as training qubit count increases. (e) Scalability assessment of GroverGPT trained on $3\sim 6$ qubits, showing accuracy (blue line) and Marked Infidelity (blue bars) when tested on larger systems ranging from $6$ to $20$ qubits, highlighting the LLM’s generalization capabilities beyond its training domain.(f) (g) (h) Model performance across a wide range of hyper-parameters (LoRA rank, batch size, learning rate, respectively), highlighting accuracy as a robust indicator.
@@ II.3 Scalability and Generalization Capability of GroverGPT
In this section, we examine GroverGPT’s remarkable ability to generalize quantum search capabilities from training on small-scale systems to significantly larger quantum systems. GroverGPT exhibits remarkable generalization capabilities across different qubit scales. Figure 3e comprehensively analyzes the model’s performance when trained on $3\sim 6$ qubits and tested on increasingly larger systems, ranging from $6$ to $20$ qubits. The results show that GroverGPT maintains accuracy above $95\%$ for systems up to $8$ qubits, with only a gradual decline to approximately $95\%$ for systems between $9$ and $20$ qubits. This robust generalization ability is particularly noteworthy given that the model was trained only on smaller systems ($3\sim 6$ qubits), suggesting its capacity to extrapolate quantum search principles to substantially larger quantum systems. The parallel trends between accuracy and infidelity in Figure 3e suggest that the model’s performance degradation at larger qubit counts is gradual and predictable, indicating a systematic rather than catastrophic breakdown of simulation capability. This behavior aligns with theoretical expectations about the scalability challenges in quantum simulation and provides valuable insights into the practical limits of classical models in capturing quantum phenomena.
@@
Through detailed fidelity analysis in this section, we investigate whether GroverGPT truly learns quantum search rather than simply approximating classical search strategies. The fidelity analysis in Figure 3d provides strong evidence that GroverGPT learns genuine quantum search rather than classical approximations. The consistently low infidelity values (below $0.005$ for $6$ qubits and approaching zero for systems trained on $5$ or more qubits) indicate that GroverGPT accurately captures the quantum state amplitudes characteristic of Grover’s algorithm. This is in stark contrast to GPT-4o openai2024gpt4ocard ’s performance, which shows a consistently high infidelity (approximately $0.011$), suggesting it fails to capture the subtle quantum features of the search process. Similarly, models such as DeepSeek-V2-Lite liu2024deepseek ; guo2025deepseek and Llama-3.2-3B vavekanand2024llama exhibit even lower accuracy and higher infidelity values, emphasizing their limitations in capturing the quantum search process. The convergence of infidelity to near-zero values with increased training qubit count is particularly significant, as it indicates that GroverGPT successfully learns the interference patterns and amplitude amplification mechanisms that are quintessential to quantum search. This distinction is further reinforced by the symmetric behavior observed in both $6$-qubit and $10$-qubit test cases, suggesting that the model has internalized fundamental quantum principles rather than memorizing specific problem instances.
@@ II.5 Impact of Prompt Design Strategies in GroverGPT
II.6 Robustness of the GroverGPT Training Strategy
In this section, we evaluate the robustness of GroverGPT by examining the limited variability among its metric indicators and investigating how diverse training sets affect its overall performance.
Ablation Study on Training Configurations. Our comprehensive ablation studies investigating LoRA rank, batch size, and learning rate, presented in Figure 3(f)(g)(h), demonstrate that GroverGPT exhibits remarkable stability across various hyper-parameter configurations. The model maintains consistent performance in terms of both accuracy and infidelity metrics under standard training conditions. However, performance degradation becomes evident at extreme parameter values, particularly when the learning rate falls below $3\times 10^{-4}$ or exceeds $1\times 10^{-3}$. These findings underscore GroverGPT’s robustness within conventional hyper-parameter ranges, suggesting a practical advantage for real-world implementations.
Figure 4: The model’s robustness across diverse training datasets on different qubit systems illustrates its sensitivity to hyper-parameters during finetuning.
Training Diversity for Robustness. Our experimental results demonstrate that training set diversity plays a crucial role in enhancing model robustness. Figure 4 illustrates this through four heatmaps generated under different hyper-parameter configurations. We systematically expanded the training dataset from a single qubit configuration (e.g., qubits=$3$) to increasingly diverse combinations (e.g., qubits=$[3,4]$, $[3,4,5]$, and $[3,4,5,6]$). The results reveal that as training data diversity increases, regions of high accuracy become more distinctly defined within the hyper-parameter space, facilitating more effective optimization. These findings suggest that enhanced data diversity not only strengthens model robustness but may also improve generalization capabilities across other challenging scenarios.
@@ III.3 Evaluation Metrics
  • Infidelity ($\epsilon$): Quantifies the overall quantum state reproduction accuracy through:
@@
  • Marked Infidelity ($\epsilon^{k}$): Specifically evaluates the accuracy of marked state probability predictions:
@@
These metrics provide complementary perspectives on GroverGPT’s performance: $\alpha$ assesses the model’s ability to identify correct search solutions, $\epsilon$ evaluates the overall quantum state reproduction infidelity, and $\epsilon^{k}$ focuses specifically on the accuracy of marked state predictions. Together, they comprehensively evaluate the model’s practical search capabilities and theoretical quantum properties.
@@ Acknowledgment
We thank Yangrui Hu, Jin-Peng Liu, Ziwen Liu, Zhiding Liang and Suqiong Zeng for discussion and help. MC, JC, and JL are supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, and the PQI Community Collaboration Awards. PL and TC are supported in part by Cisco Faculty Award and UNC SDSS Seed Grant.
@@ IV.5 GroverGPT Fine-Tuning Details
where $N$ is the number of training samples, $T_{i}$ is the length of the target sequence for the $i$-th sample, $y_{i,t}$ is the ground truth token, $y_{i,<t}$ represents the preceding tokens, $x_{i}$ is the input prompt, and $P_{\theta}(y_{i,t}|y_{i,<t},x_{i})$ is the model’s predicted probability. This loss function ensures that the model learns to predict target tokens accurately based on the provided context and previously predicted tokens.